OpenStudy (anonymous):

How do I evaluate: log(1/10)1?

OpenStudy (anonymous):

\[\log _{1/10}1\]

OpenStudy (anonymous):

the answer states: logn1=0 for all non-zero values of n

OpenStudy (anonymous):

I'm confused by that statement.. /:

OpenStudy (anonymous):

Use,\[\log_{a} x=\frac{\log_{10}x}{\log_{10} a} \]

OpenStudy (anonymous):

In this case x=1, a=1/10

OpenStudy (asnaseer):

are you aware of the rule that if:\[log_ab=c\]then:\[b=a^c\]

OpenStudy (anonymous):

yes I am

OpenStudy (asnaseer):

so if c=0, the you get:\[b=a^0=1\]

OpenStudy (anonymous):

@asnaseer yeah that works too

OpenStudy (asnaseer):

*then

OpenStudy (anonymous):

yes, ohh.. so why was there an 'n'?

OpenStudy (asnaseer):

they are just stating it in this way:\[\log_n1=0\]

OpenStudy (anonymous):

oh.. okay, thank you so much!

OpenStudy (asnaseer):

yw :)